The rate of economic growth per capita in France from 1996 to 2000 was 1.9% per year, while in Korea over the same period it was 4.2%. Per capita real GDP was $28,900 in France in 2003, and $12,700 in Korea. Assume the growth rates for each country remain the same.

Compute the doubling time for France’s per capita real GDP.
Compute the doubling time for Korea’s per capita real GDP.
What will France’s per capita real GDP be in 2045?
What will Korea’s per capita real GDP be in 2045?
1
Expert's answer
2021-08-02T00:46:03-0400

The doubling time for France’s per capita real GDP is:

n = ln(2)/ln(1.019) = 36.8 years.

The doubling time for Korea’s per capita real GDP is:

n = ln(2)/ln(1.042) = 16.8 years.

France’s per capita real GDP in 2045 will be:

"GDP = 28,900\u00d71.019^{42} = 63,710.88."

Korea’s per capita real GDP in 2045 will be:

"GDP = 12,700\u00d71.042^{42} = 71,490.43."